# MindEye-OmniAssist: A Gaze-Driven LLM-Enhanced Assistive Robot System for Implicit Intention Recognition and Task Execution

### v1

@@ I INTRODUCTION
Figure 1: The schematic diagram of the MindEye-OmniAssist. First, the user’s view of scene image and gaze point are input into the gaze-based intention recognition module we proposed. The intention recognition module first uses VFM and a custom-designed network to identify the objects the user intends to interact with, and then leverages the capabilities of the LLM to infer the user’s full intention. After the inference is complete, the system proceeds to the intention confirmation module, where it uses speech to verify whether the inferred intention is correct, and the user provides feedback via gaze. Finally, the system proceeds to the motion planning module, where it plans and executes the actions to accomplish the intended task.
Limb function plays a crucial role in enabling individuals to perform daily tasks. However, individuals with impaired limb function may struggle to independently perform tasks[1]. Assistive robotic systems can enhance and support the user’s limb capabilities[2, 3, 4]. One promising approach to facilitate effective human-robot interaction in these systems is gaze-based control[5, 6]. This approach has significant potential because the human gaze is inherently intentional and goal-directed, and individuals with limb disabilities can still maintain control over their eye movements, providing valuable spatial cues about objects with which they intend to interact. Given these advantages, gaze-based assistive systems are essential and promising.
@@
(1) A gaze-based open implicit intention recognition framework powered by LLM and VFM. This framework is capable of processing gaze input, and address undefined user intentions.
@@
Fig.1 illustrates the schematic diagram of our research. The rest of this article is organized as follows. Section II provides an overview of our system, including both the hardware and software architecture. Section III presents a detailed description of each component within the software architecture. Section IV reports the experimental setup and results. Our discussion is presented in section V. Finally, Section VI concludes this article.
@@ II-A Hardware Architecture
Figure 2: The hardware setup of MindEye-OmniAssist.
The hardware setup of the MindEye-OmniAssist consists of a Pupil Invisible head-mounted eye tracker, four collaborative robotic arms, three Intel RealSense D435 series cameras and a laptop, as shown in Fig.2. The Pupil Invisible’s scene camera is capable of outputting scene images at a resolution of $1088\times 1080$ and a rate of 30 Hz. Furthermore, the eye tracker provides realtime 2D gaze points $[g_{x},g_{y}]$ on the scene image, at a frequency of about 120 Hz. The robot has four arms, comprising two main arms and two follower arms, offering two operating modes: in teleoperation mode, the follower arms move in tandem with the main arms; in autonomous mode, the follower arms move based on code instructions, while the main arms remain stationary. The three cameras provide RGB and depth images with different fields of view. All devices communicate with a PC running Ubuntu operating system. The assistive robotic system implemented based on Robotic Operation System (ROS).
Figure 3: Open implicit intention recognition framework based on gaze and foundation models. The proposed framework employs a hierarchical reasoning mechanism. The open vocabulary object detector is first utilized to identify objects in scene images. Subsequently, an intention recognition network is applied to classify each object based on whether the user exhibits an interaction intention toward it, enabling the system to focus on objects relevant to the user’s goals. Finally, all objects associated with user interaction intentions are processed through a LLM to infer the user’s complete intentions, thereby establishing a semantic link between visual perception and user intent.
@@ III-A Gaze-based Intention Recognition
Object Detection Powered by VFM. With the rapid development of VFMs, current technologies can achieve high-precision recognition of various objects, making it possible to recognize different objects in open and complex environments. We use the open vocabulary object detector YOLO-World model [21] to identify objects in scene images and record the recognition results. For each object $o$ in the scene image, the recognition result is represented as:
@@
where $bs$ represents the number of samples input at each time, $sw$ is the length of the sliding window, and $num\_fea$ is the length of the feature vector, which in this paper is three.
Intention Recognition. The intention recognition task is a binary classification problem, where the goal is to classify each gaze input into one of two categories: “intent to interact with the object” or “no intent to interact”. We design and implement a multi-layer neural network structure as shown in Fig.4. The network consists of several key modules, including multi-scale convolutional layers, positional encoders, Transformer encoders, and fully connected layers. The input data $f$ is processed through three multi-scale convolutional layers with kernels of scales (3,7,13) to extract features from varying receptive fields [22]. Subsequently, the outputs are concatenated along the channel dimension to form a comprehensive feature representation, which is then normalized and activated using ReLU. To further improve feature selection ability, attention mechanism is introduced. The resulting convolutional features are denoted as $X_{\text{conv}}$.
To capture positional relationships within time-series data, we also introduce positional encoding to the input data $f$. The features after positional encoding are fed into a Transformer encoder [23] to model the global dependencies among the features. After processing through multiple convolution layers and Transformer encoding, the feature vector $X_{\text{trans}}$ is obtained. The information from both branches is then merged:
$X_{\text{new}}=\text{concat}(X_{\text{conv}},X_{\text{trans}}).$ (5)
The merged features are mapped to a new feature space through a fully connected layer, and a Sigmoid activation function is used to produce the final classification result $\hat{y}$. If $\hat{y}$ is greater than 0.5, it means there is an intention to interact with the object; otherwise, there is no intention to interact.
Intention Inference Based on LLM. Implicit intention recognition requires inferring the user’s complete intentions based on identifying which objects they intend to interact with. Models trained on large corpora possess a robust knowledge base and inferential capabilities. Utilizing language models to understand common object combinations and their associated activities allows for an effective integration of low-level object recognition with high-level behavioral inference. We leverage the knowledge base of DeepSeek-R1 [24] to deduce users’ intentions in various scenarios, as illustrated in Fig.5.
Figure 5: An example of implicit intention inference. When a user intends to interact with coffee, cup, and kettle, the intention could be to drink coffee or make coffee.
@@ III-B Intention Confirmation
Once the user’s intention $\langle intention\rangle$ is inferred, the system will confirm it with the user via voice: “Is your intention $\langle intention\rangle$?”. Upon hearing this confirmation, if the inferred intention is correct, the user is expected to look at Area 3 to indicate agreement. Conversely, if the intention is incorrect, the user should look at Area 1 to reject the execution of the inferred intention. This method ensures that the system receives explicit feedback from the user before proceeding with any action, improving the accuracy and reliability of the system.
@@ III-C Motion Planning
For example, when the user has an interactive intention with a pear and a fruit bowl, the system infers the intention as “place the pear into the fruit bowl.” After the user confirms this intention, the LLM generates the action sequence: first, locate the pear and the fruit bowl, then grasp the pear and place it into the fruit bowl. The generated action sequence is used to call the corresponding APIs, guiding the robotic arm’s movement.
@@ IV-A Experimental Setup
Figure 7: The experimental setup of data collection.
Figure 8: A subject is performing the task of watering the plants with the assistance of the robotic arm. The figure illustrates four perspectives, through which one can observe the process of the user expressing the intention, confirming the intention, and the robotic arm performing the task.
Dataset. To complete the above evaluation, we need a dataset where users express implicit intentions through eye movements. Since there is no such open-source dataset available at present, we established an eye movement dataset using the Pupil Invisible eye tracker. The dataset includes data from eight volunteers, aged 20 to 24, and contains various objects, including bottles, scissors, bananas, grapes, etc. The experimental setup involved subjects sitting in front of a table with these objects, wearing the eye tracker, and performing tasks designed to express interaction intentions through gaze, as shown in Fig.7. In one part of the experiment, participants expressed their intention to interact with objects, such as gazing at a banana to signal picking it up, or gazing at a kettle and cup to indicate pouring water. In another part, participants gazed unconsciously to simulate natural eye movement. The position of objects was changed throughout the experiment and each task was repeated five times. Eye tracker calibration was performed before each round to ensure accurate gaze estimation, and recalibration occurred if accuracy dropped during the experiment. Approximately 120 minutes of data were collected across all trials.
After data collection, several post-processing steps were applied to align the data along the time axis and mitigate the influence of head movements on gaze points. Additionally, the intentions behind participants’ eye movements were manually annotated to create labeled data for training and evaluation.
2) System Evaluation: We tested our system in a desktop environment. Five participants took part in the system evaluation experiment. In the experiment, subjects wore the eye-tracker and were seated approximately 35 cm away from the table. The eye-tracker was calibrated prior to the start of each trial to ensure accurate gaze tracking. If any drift or inaccuracy was observed during the experiment, recalibration was performed to maintain data precision. During the experiment, the participants’ gaze was not disturbed, allowing for natural behaviors, such as blinking.
@@
During each trial, 1 to 5 different objects were randomly placed on the table. These objects included items such as bottles, cups, bowls, etc. The participants were not given specific instructions regarding the tasks; instead, they were free to express their intentions through gaze, and the robotic arm would assist in executing the tasks accordingly. A trial was considered successful if the task was completed within three attempts. For each trial, the success rate at different stages is recorded at the same time. The performance of the system was assessed in terms of the overall success rate, the success rate of intention recognition, the success rate of action planning and the success rate of action execution. Throughout the experiment, we recorded the reason for any failures. Fig.8 shows a subject completing the task of watering the plants with the assistance of the robotic arm.
@@ IV-B Experimental Results and Discussion
The MindEye-OmniAssist system achieved good performance across all tasks, with particularly high success rates in the recognition and planning stages. For tasks like fetching an object and putting something into something, the system achieved perfect recognition (20/20 and 18/20) and planning (20/20 and 17/18), with slightly lower execution success (16/20 and 15/17). Watering the plants showed perfect recognition and planning (5/5), but lower execution (3/5). In toggling the switch, recognition was good (4/5), while planning (3/4) and execution (3/3) had minor issues. For pouring water, the system performed well with perfect recognition and planning (5/5), and a slightly lower execution success (4/5). Overall, the system achieved an overall success rate of 41/55, with high recognition (52/55) and planning (50/52) success rates. The execution success rate was 41/50, highlighting that while the system performs well in recognizing user intentions and planning actions, execution remains a challenge, especially in tasks requiring more complex manipulations.
V Discussion
The MindEye-OmniAssist system exhibited strong performance in intention recognition and action planning across various tasks, though execution success showed some variability. Compared to prior studies, this system stands out for its ability to handle a diverse range of assistive tasks, all driven by gaze input. It successfully demonstrates the potential for a gaze-based interface to recognize complex user intentions and adapt to a broader set of tasks.
Although the results above show some positive signs, the system is not perfect. First, the main reasons for task failures in the experiment were failures in the robotic arm action execution, such as failure to avoid obstacles during execution, and insufficient execution accuracy. The main reason for this is issues with grounding operations. Current system only plans actions based on user intent, without real-time information about the environment. This limits the system’s ability to handle action details and avoid obstacles, which lowers the success rate. In the future, work can focus on improving the system’s ability to sense and integrate environmental information to enhance action planning and obstacle avoidance. Additionally, the basic actions of the system are learned through behavior cloning, so the action set is limited. When more basic actions are needed, the system currently relies on manual teaching and behavior cloning to expand its action library. Therefore, future work could consider using automated learning or more efficient methods to update and expand the basic action library, making the system more flexible and adaptable.
VI CONCLUSIONS
Gaze-based control in assistive robotic systems holds promise for enhancing upper limb capabilities, but the limited intent recognition restricts current systems to basic tasks. In this study, we proposes a gaze-based open implicit intention recognition framework based on LLM and VFM, capable of recognizing user intentions beyond predefined scenarios. Specifically, an open-vocabulary object detector is used to identify objects in the scene, and an intention recognition network classifies each object based on whether it is a target for interaction. Interaction targets are then input into the LLM to infer the user’s complete intention. In addition, we present a gaze-driven assistive robotic system. After inferring the intention through gaze signals, the system provides the intention to the user through voice. The user provides corrective feedback through eye movement signals, and the system subsequently generates customized action sequences based on the LLM to assist the user in completing tasks. The experimental results indicate that our system can effectively identify user intentions and assist users in completing various tasks. Future work will involve improving the accuracy of action planning and developing more effective methods for updating the basic action library.

### v2

@@ I INTRODUCTION
Figure 1: The schematic diagram of the MindEye-OmniAssist.
Limb function plays a crucial role in enabling individuals to perform daily tasks. However, individuals with impaired limb function may struggle to independently perform tasks[1]. Assistive robotic systems can enhance and support the user’s limb capabilities[2]. One promising approach to facilitate effective human-robot interaction in these systems is gaze-based control[3, 4]. This approach has significant potential because the human gaze is inherently intentional and goal-directed, and individuals with limb disabilities can still maintain control over their eye movements, providing valuable spatial cues about objects with which they intend to interact. Given these advantages, gaze-based assistive systems are essential and promising.
@@
(1) A gaze-based open implicit intention recognition framework powered by LLM and VFM. This framework is capable of processing gaze input, and addressing undefined user intentions.
@@ II-A Hardware Architecture
The hardware setup of the MindEye-OmniAssist consists of a Pupil Invisible head-mounted eye tracker, four collaborative robotic arms, three Intel RealSense D435 series cameras and a laptop . The Pupil Invisible’s scene camera is capable of outputting scene images at a resolution of $1088\times 1080$ and a rate of 30 Hz. Furthermore, the eye tracker provides realtime 2D gaze points $g_{x},g_{y}]$ on the scene image, at a frequency of about 120 Hz. The robot has four arms, comprising two main arms and two follower arms, offering two operating modes: in teleoperation mode, the follower arms move in tandem with the main arms; in autonomous mode, the follower arms move based on code instructions, while the main arms remain stationary. The three cameras provide RGB and depth images with different fields of view. All devices communicate with a PC running Ubuntu operating system. The assistive robotic system implemented based on Robotic Operation System (ROS).
Figure 2: Open implicit intention recognition framework based on gaze and foundation models.
@@ III-A Gaze-based Intention Recognition
Object Detection Powered by VFM. We use the open vocabulary object detector YOLO-World model [19] to identify objects in scene images and record the recognition results. For each object $o$ in the scene image, the recognition result is represented as:
@@
where $bs$ represents the number of samples input at each time, $sw$ is the length of the sliding window, and num_feanum\_fea is the length of the feature vector, which in this paper is three.
Intention Recognition. The network consists of several key modules, including multi-scale convolutional layers, positional encoders, Transformer encoders, and fully connected layers. The input data $f$ is processed through three multi-scale convolutional layers with kernels of scales (3,7,13) to extract features from varying receptive fields [20]. Subsequently, the outputs are concatenated along the channel dimension to form a comprehensive feature representation. To further improve feature selection ability, attention mechanism is introduced. The resulting convolutional features are denoted as $X_{\text{conv}}$.
To capture positional relationships within time-series data, we also introduce positional encoding to the input data $f$. The features after positional encoding are fed into a Transformer encoder [21] to model the global dependencies among the features. After the feature vector $X_{\text{trans}}$ is obtained, the information from both branches is then merged and mapped to a new feature space through a fully connected layer, and a Sigmoid activation function is used to produce the final classification result $\hat{y}$. If $\hat{y}$ is greater than 0.5, it means there is an intention to interact with the object; otherwise, there is no intention to interact.
Intention Inference Based on LLM. Implicit intention recognition requires inferring the user’s complete intentions based on identifying which objects they intend to interact with. Models trained on large corpora possess a robust knowledge base and inferential capabilities. Utilizing language models to understand common object combinations and their associated activities allows for an effective integration of low-level object recognition with high-level behavioral inference. We leverage the knowledge base of DeepSeek-R1 [22] to deduce users’ intentions in various scenarios .
@@ III-B Intention Confirmation
Once the user’s intention $\langle intention\rangle$ is inferred, the system will confirm it with the user via voice: “Is your intention $\langle intention\rangle$?”. Upon hearing this confirmation, if the inferred intention is correct, the user is expected to look at Area 3 to indicate agreement. Conversely, if the intention is incorrect, the user should look at Area 1 to reject the execution of the inferred intention.
@@ IV-A Experimental Setup
Dataset. To complete the above evaluation, we need a dataset where users express implicit intentions through eye movements. Since there is no such open-source dataset available at present, we established an eye movement dataset using the Pupil Invisible eye tracker. The dataset includes data from eight volunteers, aged 20 to 24, and contains various objects, including bottles, scissors, bananas, grapes, etc. The experimental setup involved subjects sitting in front of a table with these objects, wearing the eye tracker, and performing tasks designed to express interaction intentions through gaze . In one part of the experiment, participants expressed their intention to interact with objects, such as gazing at a banana to signal picking it up, or gazing at a kettle and cup to indicate pouring water. In another part, participants gazed unconsciously to simulate natural eye movement. The position of objects was changed throughout the experiment and each task was repeated five times. Eye tracker calibration was performed before each round to ensure accurate gaze estimation, and recalibration occurred if accuracy dropped during the experiment. Approximately 120 minutes of data were collected across all trials.
After data collection, several post-processing steps were applied to align the data along the time axis and mitigate the influence of head movements on gaze points. Additionally, the intentions behind participants’ eye movements were manually annotated to create labeled data for training and evaluation. Therefore, data at each moment in the dataset includes a scene image, 2D gaze point coordinates relative to the scene image, and manually annotated labels indicating the presence of interaction intent.
2) System Evaluation: We tested our system in a desktop environment. Five participants took part in the experiment.Subjects wore the eye-tracker and were seated approximately 35 cm away from the table. The eye-tracker was calibrated prior to the start of each trial to ensure accurate gaze tracking. If any drift or inaccuracy was observed during the experiment, recalibration was performed to maintain data precision. During the experiment, the participants’ gaze was not disturbed, allowing for natural behaviors, such as blinking.
@@
During each trial, 1 to 5 different objects were randomly placed on the table. These objects included items such as bottles, cups, bowls, etc. The participants were not given specific instructions regarding the tasks; instead, they were free to express their intentions through gaze, and the robotic arm would assist in executing the tasks accordingly. A trial was considered successful if the task was completed within three attempts. For each trial, the success rate at different stages is recorded at the same time. The performance of the system was assessed in terms of the overall success rate, the success rate of intention recognition, the success rate of action planning and the success rate of action execution. Throughout the experiment, we recorded the reason for any failures.
@@ IV-B Experimental Results and Discussion
V CONCLUSIONS
Gaze-based control in assistive robotic systems holds promise for enhancing upper limb capabilities, but the limited intent recognition restricts current systems to basic tasks. In this study, we proposes a gaze-based open implicit intention recognition framework based on LLM and VFM, capable of recognizing user intentions beyond predefined scenarios. Specifically, an open-vocabulary object detector is used to identify objects in the scene, and an intention recognition network classifies each object based on whether it is a target for interaction. Interaction targets are then input into the LLM to infer the user’s complete intention. In addition, we present a gaze-driven assistive robotic system. After inferring the intention through gaze signals, the system provides the intention to the user through voice. The user provides corrective feedback through eye movement signals, and the system subsequently generates customized action sequences based on the LLM to assist the user in completing tasks. The experimental results indicate that our system can effectively identify user intentions and assist users in completing various tasks. Future work will focus on improving the accuracy of action planning , developing more effective methods for updating the basic action library, and implementing error correction mechanisms.